If the roots of the given equation 2x^2+3(A-2)x+ A+4=0 be equal in magnitude but opposite in sign, then A=
(a)1
(b)2/3
(c)3
(d)2

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
hence if one root is t and other root will be – t
 
hence sum of oots = t + (- t) = – 3(A – 2)/2
 
- 3(A- 2) /2 = 0
 
A – 2 = 0
 
A = 2
 
Jope it helps